OpenStudy (anonymous):

find the quotient: 25c^4+20c^3 ------------ 5c

OpenStudy (anonymous):

when you divide with the same base you subtract the exponents 5c-c has an exponent of 1

OpenStudy (anonymous):

25c^4/5c^1=5c^3 and 20c^3/5c^1=4c^2

OpenStudy (anonymous):

so final answer is 5c^3+4c^2...do you understand?

OpenStudy (anonymous):

yea thanks

OpenStudy (anonymous):

you're welcome...anything else?

OpenStudy (anonymous):

yea is this right 63b^4-9b^3 --------- 9b = -7b^3+b^2

OpenStudy (anonymous):

yes or 7b^3-b^2

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

thanks for ur help

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

-21w^4-21w^3 ------------- 7w = -3w^3-3w^2 is this right?

OpenStudy (anonymous):

i thnk i messed up it should be a + not - right?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

you are correct!

OpenStudy (anonymous):

u no much bout inequalities

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

y<x=2 \[y \ge5\]

OpenStudy (anonymous):

y<x+2 sorry

OpenStudy (anonymous):

n second one together

OpenStudy (anonymous):

have to graph it too

OpenStudy (anonymous):

do you know how to graph using y=mx+b with y-int and slope

OpenStudy (anonymous):

n second one together

OpenStudy (anonymous):

yea kind of confusing

OpenStudy (anonymous):

ok second inequality is a horizontal line passing throug (0, 5) shading above the solid line

OpenStudy (anonymous):

first one is a dotted line using y=mx+b the y-int(b=2) you start at (0,2) and use the slope(m-coefficient of x) which is 1/1 so from (0,2) rise 1(up) run 1(right) to get second point

OpenStudy (anonymous):

shading should include (0,0) so to the right of the diagonal line

OpenStudy (anonymous):

the solution is the area where the shading overlaps-top right area similar to less than shape : )

OpenStudy (anonymous):

did you get lost?

OpenStudy (anonymous):

yeah does it look like this kind of lol the direction \ ---------------- \

OpenStudy (anonymous):

the diagonal lines is going up from left to right bc the slope is positive

OpenStudy (anonymous):

like this lol / --------------------- / /

OpenStudy (anonymous):

so lower right is shaded right

OpenStudy (anonymous):

top right---see pic but diagonal should be dotted

OpenStudy (anonymous):

k thas what i have but had top line dotted too lol appreciate ur patience:)

OpenStudy (anonymous):

dotted line for greater than or less than solid when you have equal sign with the inequality

OpenStudy (anonymous):

love helping students in math so I have great patience : )

OpenStudy (anonymous):

y>4 \[x \le2\] up two straight across on fourth line

OpenStudy (anonymous):

2 is diagonal right dotted line

OpenStudy (anonymous):

y>4 is dotted line through (0,4) shading above the line because when you substitute (0,0) into the inequality, 0>4 is not true so shading should not include test point (0,0)

OpenStudy (anonymous):

2nd one is vertical solid line passing through (2,0) shade to the left because 0 is less than or equal to 2

OpenStudy (anonymous):

solution area is in top left right angle

OpenStudy (anonymous):

I think i got it right i would post pic but when other viewer clicks it opens blank dont no why.

OpenStudy (anonymous):

what is your number i can send you a text with pic

OpenStudy (anonymous):

nevermind I'll send to my email and post in here

OpenStudy (anonymous):

but anyways thank you so much.... had to take a plato class fot algebra. i graduate may 26, this month so this will be my last open study. thanks again

OpenStudy (anonymous):

CONGRATS!!! Do you need me to post pic

OpenStudy (anonymous):

yea please

OpenStudy (anonymous):

rsanchz68@gmail.com

OpenStudy (anonymous):

you can email me if you need any other math help else brishundra@yahoo.com

OpenStudy (anonymous):

where is diagonal line at

OpenStudy (anonymous):

no diagonal line is this one becuase y= is horizontal x=is vertical

OpenStudy (anonymous):

so same is true for inequality graphs..graph them just like y= or x=

OpenStudy (anonymous):

o yeah k got ur email thanks

OpenStudy (anonymous):

welcome!